### v1

# On denominators of consecutive $\operatorname{SL}(2,\mathbb{N})$-saturated Farey fractions

Abstract.
The sequences of $\operatorname{SL}(2,\mathbb{N})$-saturated Farey fractions were defined in our previous work as ${\mathscr{S}}_{Q}:=\{a/q\in{\mathbb{Q}}\cap(0,1]:q+a+\bar{a}\leq Q\}$, where $\bar{a}$ is the multiplicative inverse of $a\left(\mathrm{mod\;}q\right)$ in $[1,q)$. Here, we prove that the set of $Q$-scaled denominators of consecutive fractions in ${\mathscr{S}}_{Q}$ is dense in the region ${\mathcal{V}}:=\{(x,y)\in[0,1]^{2}:\max\{(1-3x)/2,2x-1\}\leq y\leq\max\{x,1-x\}\}$, and provide a formula for their distribution in $\mathcal{V}$} as $Q\rightarrow\infty$.
2020 Mathematics Subject Classification:
Primary 11B57. Secondary: 11J71, 11K36, 11L05.
Key words and phrases: $\operatorname{SL}(2,{\mathbb{N}})$-saturated Farey fractions; mediant insertion; consecutive denominators
1. Introduction
Motivated by the study of the distribution of reduced quadratic irrationals [5, 12, 19, 21, 22], we considered in [3] the sequence $(\mathscr{S}_{Q})_{Q}$ of $\operatorname{SL}(2,\mathbb{N})$-saturated Farey fractions, defined by
$\mathscr{S}_{Q}:=\left\{\frac{d}{b}:\begin{matrix}\text{ there exist }a,c\in\mathbb{N}\text{ such that }\left(\begin{smallmatrix}a&b\\ c&d\end{smallmatrix}\right)\in\operatorname{SL}(2,\mathbb{Z})\\ a\geq b\geq d\geq 1,\ a\geq c\geq d,\ a+d\leq Q\end{matrix}\right\},\quad Q\geq 3,$
viewed as a subset of the customary Farey set
$\mathcal{F}_{Q}:=\left\{\frac{d}{b}:d,b\in\mathbb{N},\ 1\leq d\leq b\leq Q,\ (d,b)=1\right\}.$
As observed in [3], this set can also be conveniently described as
$\mathscr{S}_{Q}=\bigg{\{}\frac{a}{q}\in\mathbb{Q}\cap(0,1]:h\bigg{(}\frac{a}{q}\bigg{)}:=q+a+\bar{a}\leq Q\bigg{\}},\quad Q\geq 3,$
where $\bar{a}$ denotes the multiplicative inverse of $a\left(\mathrm{mod\;}q\right)$ in $[1,q)$. In particular, this shows that $\mathcal{F}_{Q}\subseteq\mathscr{S}_{3Q}$. Note also that $h(a/q)\neq h(1-a/q)$, but some sort of $(x\mapsto 1-x)$-symmetry of the height $h$ is present in the identity
$h\bigg{(}\frac{a}{q}\bigg{)}+h\bigg{(}1-\frac{a}{q}\bigg{)}=4q.$
The sets $\mathscr{S}_{Q}$ exhibit interesting arithmetic properties, resembling in part those of $\mathcal{F}_{Q}$. Firstly, their union gives precisely the set ${\mathbb{Q}}\cap(0,1]$. Secondly, as shown in the appendix to [3], the elements of each set $\mathscr{S}_{Q}^{*}:=\{0\}\cup\mathscr{S}_{Q}$ define a modular partition of the interval $[0,1]$. This means that if $0/1<a_{1}/q_{1}=1/(Q-2)<\cdots<a_{\#\mathscr{S}_{Q}}/b_{\#\mathscr{S}_{Q}}=1/1$ denotes the elements of $\{0\}\cup\mathscr{S}_{Q}$, then $a_{i+1}q_{i}-a_{i}q_{i+1}=1$ for any $i=1,\ldots,\#\mathscr{S}_{Q}-1$.
In sharp contrast with the sequence $(\mathcal{F}_{Q})_{Q}$, the elements of $(\mathscr{S}_{Q})_{Q}$ are no longer uniformly distributed as $Q\rightarrow\infty$. Furthermore, it was proved in [3, Theorem 1] that
$\big{(}\mathscr{S}_{Q}\cap[0,\beta]\big{)}=\frac{Q^{2}}{2\zeta(2)}\log\bigg{(}\frac{2+2\beta}{2+\beta}\bigg{)}+O_{\varepsilon}(Q^{3/2+\varepsilon}),\quad\forall\beta\in[0,1].$
In particular, we have
$\lim\limits_{Q}\frac{\#(\mathscr{S}_{Q}\cap[0,1/2])}{\#\mathscr{S}_{Q}}=\frac{\log(6/5)}{\log(4/3)}\approx 0.63376.$ (1)
Figure 1. The set of pairs of denominators of consecutive fractions in $\mathscr{S}_{1200}$. The points are colored differently, depending on the number of intermediate Farey fractions whose insertion is delayed at this level.
One of the fundamental features of the Farey sequence $(\mathcal{F}_{Q})_{Q}$ is that the elements of the set $\mathcal{F}^{*}_{Q}:=\{0\}\cup\mathcal{F}_{Q}$ form a modular partition of the interval $[0,1]$. In particular, the elements of $\mathcal{F}^{*}_{Q}$ are determined by their denominators. The elements of the set
$\mathcal{D}_{Q}:=\big{\{}(q_{1},q_{2}):\exists\,a_{1},a_{2}\in\mathbb{Z},\ a_{1}/q_{1}<a_{2}/q_{2}\text{ consecutive elements in }\mathcal{F}^{*}_{Q}\big{\}}$
are in one-to-one correspondence with the set of primitive lattice points in the triangle $Q\mathcal{T}$, where $\mathcal{T}$ denotes the Farey triangle $\{(x,y)\in[0,1]^{2}:x+y>1\}$. As a result, the scaled set $\mathcal{D}:=\cup_{Q}Q^{-1}\mathcal{D}_{Q}$ is dense in the closed triangle $\overline{\mathcal{T}}$. It was first noticed by Kargaev and Zhigljavsky [13] (see also [6, Lemma 2] for a quick effective proof) that the $Q$-scaled sets $Q^{-1}\mathcal{D}_{Q}$ are uniformly distributed in $\overline{\mathcal{T}}$, in the sense that, for every rectangular box $\mathcal{R}\subseteq\overline{\mathcal{T}}$,
$\lim\limits_{Q}\frac{\#(\mathcal{D}_{Q}\cap Q\mathcal{R})}{\#\mathcal{D}_{Q}}=\operatorname{Area}(\mathcal{R}).$ (2)
As noticed in [6], the shifting of consecutive elements in $\mathcal{D}_{Q}$ is conveniently described by iterating the map $T:\mathcal{T}\rightarrow\mathcal{T}$,
$T(x,y):=\big{(}y,\kappa(x,y)y-x\big{)}\quad\text{with }\kappa(x,y):=\bigg{\lfloor}\frac{1+x}{y}\bigg{\rfloor},$
which satisfies, for every $Q\geq 1$,
$T\bigg{(}\frac{q_{1}}{Q},\frac{q_{2}}{Q}\bigg{)}=\bigg{(}\frac{q_{2}}{Q},\frac{q_{3}}{Q}\bigg{)},$
if $a_{1}/q_{1}<a_{2}/q_{2}<a_{3}/q_{3}$ are consecutive elements in $\mathcal{F}^{*}_{Q}$.
The primary aim of this paper is to investigate the existence of an analogue of (2) when $\mathcal{F}_{Q}$ is being replaced by $\mathscr{S}_{Q}$. For every $Q\geq 3$, we consider the set $\mathscr{D}_{Q}$ of denominators $(q_{1},q_{2})$ of consecutive elements $a_{1}/q_{1}<a_{2}/q_{2}$ from $\mathscr{S}_{Q}$. The role of $\overline{\mathcal{T}}$ will be played here by the region
$\mathcal{V}:=\bigg{\{}(x,y)\in[0,1]^{2}:\max\bigg{\{}\frac{1-3x}{2},2x-1\bigg{\}}\leq y\leq\max\{1-x,x\}\bigg{\}}.$
Theorem 1.
The set $\mathscr{D}:=\cup_{Q}Q^{-1}\mathscr{D}_{Q}$ is dense in $\mathcal{V}$.
Theorem 2.
The $Q$-scaled sets $Q^{-1}\mathscr{D}_{Q}$ have an asymptotic limiting distribution, described as follows. Partition the set $\mathcal{V}$ as $\mathcal{V}_{1}\cup\mathcal{V}_{2}\cup\mathcal{V}_{3}$, with
$\begin{split}&\mathcal{V}_{1}:=\bigg{\{}(x,y)\in\bigg{[}\frac{1}{2},1\bigg{]}\times\bigg{[}\frac{1}{3},1\bigg{]}:\max\{1-x,2x-1\}\leq y\leq x\bigg{\}},\\ &\mathcal{V}_{2}:=\bigg{\{}(x,y)\in\bigg{[}0,\frac{1}{2}\bigg{]}\times\bigg{[}\frac{1}{5},1\bigg{]}:\max\bigg{\{}\frac{1-3x}{2},x\bigg{\}}\leq y\leq 1-x\bigg{\}},\\ &\mathcal{V}_{3}:=\bigg{\{}(x,y)\in\bigg{[}\frac{1}{5},\frac{2}{3}\bigg{]}\times\bigg{[}0,\frac{1}{2}\bigg{]}:\max\bigg{\{}\frac{1-3x}{2},0,2x-1\bigg{\}}\leq y\leq\min\{x,1-x\}\bigg{\}}.\end{split}$ (3)
For every rectangular box $\mathcal{R}\subseteq\mathcal{V}_{i}$, $i=1,2,3$,
$\lim\limits_{Q}\frac{\#(\mathscr{D}_{Q}\cap Q\mathcal{R})}{\#\mathscr{D}_{Q}}=\iint_{\mathcal{R}}H_{i}(x,y)\,dxdy,$
with $H_{i}(x,y)>0$ on $\mathring{\mathcal{V}}_{i}$ given by (8), (11) and (13).
Figure 2. The polygons defined in (3) and the partition of the $\mathcal{V}$-shape described in Section 4. The order in which the polygons are positioned is from right to left and from top to bottom.
The region $\mathcal{V}$ can be partitioned into subregions $\mathcal{W}_{1},\mathcal{W}_{2},\mathcal{W}_{3},\ldots$, with $\mathcal{W}_{r}$ representing the closure of the scaled set $\cup_{Q}Q^{-1}\mathscr{D}_{Q}^{(r)}$, where $\mathscr{D}_{Q}^{(r)}$ denotes the set of pairs of denominators of consecutive elements of $\mathscr{S}_{Q}$ with exactly r1r-1 elements of $\mathcal{F}_{Q}$ in between (see the representation in the right side of Figure 2). In addition, every $\mathcal{W}_{r}$, $r\geq 3$, can be further partitioned into r1r-1 (possibly empty) polygons $\mathcal{W}_{r,i}$, with $i$ being the (unique) place where the mediant of consecutive denominators is being attained, that is $q_{i}=q_{1}+q_{r+1}$. The boundaries of the polygons $\mathcal{W}_{r,i}$ are computable, as illustrated in Section 4 for $r\leq 5$. See also Figures 1 and 2.
If the rational numbers $a_{1}/q_{1}<a_{2}/q_{2}<a_{3}/q_{3}$ satisfy $a_{2}q_{1}-a_{1}q_{2}=1=a_{3}q_{2}-a_{2}q_{3}$, then
$\frac{q_{1}+q_{3}}{q_{2}}=\frac{a_{1}+a_{3}}{a_{2}}=a_{3}q_{1}-a_{1}q_{3}\in\mathbb{N}.$
Furthermore, when $a_{1}/q_{1}<a_{2}/q_{2}<a_{3}/q_{3}$ are consecutive in $\mathcal{F}_{Q}$, we have
$\frac{q_{1}+q_{3}}{q_{2}}=\bigg{\lfloor}\frac{Q+q_{1}}{q_{2}}\bigg{\rfloor}=\kappa\bigg{(}\frac{q_{1}}{Q},\frac{q_{2}}{Q}\bigg{)}=:\nu\bigg{(}\frac{a_{2}}{q_{2}}\bigg{)}.$
This quantity, called the index of the fraction $a_{2}/q_{2}$ (in $\mathcal{F}_{Q}$), plays a critical role in the distribution of the Farey sequence $(\mathcal{F}_{Q})_{Q}$ and has been investigated by several authors (see, e.g., [6, 10, 7, 2, 1, 8, 14, 15, 17, 16]). It would be interesting to find an analogue formula for the index $(q_{1}+q_{3})/q_{2}$ when $a_{1}/q_{1}<a_{2}/q_{2}<a_{3}/q_{3}$ are consecutive elements in $\mathscr{S}_{Q}$.
One nice feature of the index, proved in [10, Theorem 1] through an elementary mediant argument, is the identity
$\sum\limits_{i=1}^{R_{Q}}\nu(f_{i})=3R_{Q}-1,$ (4)
where $R_{Q}:=\#\mathcal{F}_{Q}=\varphi(1)+\varphi(2)+\cdots+\varphi(Q)$, $f_{1}=1/Q<\cdots<f_{R_{Q}}=1/1$ denote the elements of $\mathcal{F}_{Q}$, $f_{0}:=0/1$ and $f_{1+R(Q)}:=f_{1}$. In Section 5 we show that a similar identity holds when $\mathcal{F}_{Q}$ is replaced by $\mathscr{S}_{Q}$. Setting $S_{Q}:=\#\mathscr{S}_{Q}$, we enumerate the elements of $\mathscr{S}_{Q}$ as $\gamma_{1}=1/(Q-2)<\gamma_{2}<\cdots<\gamma_{S_{Q}}=1/1$, and take $\gamma_{0}:=0/1$, $\gamma_{1+S_{Q}}:=\gamma_{1}$, $\gamma_{i}=a_{i}/q_{i}$. A quick argument as in the proof of (4) provides the following statement.
Proposition 3.
For every $Q\geq 3$, we have
$\sum\limits_{i=1}^{S_{Q}}\frac{q_{i-1}+q_{i+1}}{q_{i}}=3S_{Q}-1.$
Figure 3. The highly oscillatory behavior of $\Phi(Q)$.
Here, the function
$\Phi(Q):=S_{Q}-S_{Q-1}=\#\big{\{}a/q\in\mathbb{Q}\cap(0,1]:q+a+\bar{a}=Q\big{\}}$
arises as a substitute for the Euler function $\varphi(Q)=R_{Q}-R_{Q-1}$, measuring the number of mediant insertions from level Q1Q-1 to level $Q$. For the correlated study of pairs $(a/q,\bar{a}/q)$ see [11, 20, 4] and the references therein. Nevertheless, the arithmetic function $\Phi$ is not multiplicative and exhibits an erratic behavior (see Figure 3).
The asymmetry of the fractions in $\mathscr{S}_{Q}$, noticed in (1), does not imply that in the case of symmetric fractions the one on the left will always appear before the one on the right by insertion. For example, $5/9\in\mathscr{S}_{16}$, but $4/9$ first appears in $\mathscr{S}_{20}$. On a different note, we remark that the circle packing generated by $\mathscr{S}_{Q}$ belongs to the class of those produced by various rules that delay the insertion (see the examples in [18, 9]), but ultimately, even though the partial configurations are different, the integer packing produced by $\mathscr{S}_{Q}$ will coincide with that produced by $\mathcal{F}_{Q}$.
2. The inclusion $\mathscr{D}\subseteq\mathcal{V}$
Let $r\geq 1$. Suppose that $a_{1}/q_{1}<\cdots<a_{r+1}/q_{r+1}$ are consecutive elements in $\mathcal{F}_{Q}$ with $a_{1}/q_{1},a_{r+1}/q_{r+1}\in\mathscr{S}_{Q}$ and $a_{2}/q_{2},\ldots,a_{r}/q_{r}\notin\mathscr{S}_{Q}$ (in other words $a_{1}/q_{1}<a_{r+1}/q_{r+1}$ are consecutive elements in $\mathscr{S}_{Q}$). Denote by $\mathscr{D}_{Q}^{(r)}$ the set of denominator pairs $(q_{1},q_{r+1})$ of consecutive elements of $\mathscr{S}_{Q}$, with $q_{1}$ and $q_{r+1}$ as above.
We start with the easiest case $r=1$.
Lemma 4.
$\mathscr{D}_{Q}^{(1)}\subseteq Q\mathcal{V}_{1}$.
Proof.
As in Section 3.1 of [3] and employing the notation thereby, in this case we have $q_{2}<q_{1}$, so $q_{1}\in[Q/2,Q]$, and also $3q_{1}-Q-q_{2}\leq\bar{q}_{2}<q_{1}$, leading to $q_{2}>2q_{1}-Q$. We also employ the trivial inequality $q_{2}>Q-q_{1}$. ∎
The situation $r\geq 2$ requires additional work. The following critical equality was proved in the appendix to [3]:
$a_{r+1}q_{1}-a_{1}q_{r+1}=1.$ (5)
We consider the index of $a_{i}/q_{i}$ in $\mathcal{F}_{Q}$, given by
$\nu_{i}:=\kappa\bigg{(}\frac{q_{i-1}}{Q},\frac{q_{i}}{Q}\bigg{)}=\bigg{\lfloor}\frac{Q+q_{i-1}}{q_{i}}\bigg{\rfloor},$
so that $q_{i+1}=\nu_{i}q_{i}-q_{i-1}$, or equivalently
$(q_{i+1},q_{i})=(q_{i},q_{i-1})\left(\begin{matrix}\nu_{i}&1\\ -1&0\end{matrix}\right),\quad i=2,\ldots,r.$
The noncommutative polynomials defined by $K^{F}_{-1}(\cdot)=0$, $K^{F}_{0}(\cdot)=1$, $K^{F}_{1}(x)=x$, and
$K^{F}_{\ell}(x_{1},\ldots,x_{\ell})=x_{\ell}K^{F}_{\ell-1}(x_{1},\ldots,x_{\ell-1})-K^{F}_{\ell-2}(x_{1},\ldots,x_{\ell-2}),$
satisfy
$\left(\begin{matrix}K^{F}_{i-1}(\nu_{2},\ldots,\nu_{i})&K^{F}_{i-2}(\nu_{2},\ldots,\nu_{i-1})\\ -K^{F}_{i-2}(\nu_{3},\ldots,\nu_{i})&-K^{F}_{i-3}(\nu_{3},\ldots,\nu_{i-1})\end{matrix}\right)=\left(\begin{matrix}\nu_{2}&1\\ -1&0\end{matrix}\right)\cdots\left(\begin{matrix}\nu_{i}&1\\ -1&0\end{matrix}\right),\quad i\geq 2,$ (6)
leading to
$(q_{i+1},q_{i})=(q_{2},q_{1})\left(\begin{matrix}K^{F}_{i-1}(\nu_{2},\ldots,\nu_{i})&K^{F}_{i-2}(\nu_{2},\ldots,\nu_{i-1})\\ -K^{F}_{i-2}(\nu_{3},\ldots,\nu_{i})&-K^{F}_{i-3}(\nu_{3},\ldots,\nu_{i-1})\end{matrix}\right),\quad i=2,\ldots,r.$
Since $a_{i+1}=\nu_{i}a_{i}-a_{i-1}$, we also have
$(a_{i+1},a_{i})=(a_{2},a_{1})\left(\begin{matrix}K^{F}_{i-1}(\nu_{2},\ldots,\nu_{i})&K^{F}_{i-2}(\nu_{2},\ldots,\nu_{i-1})\\ -K^{F}_{i-2}(\nu_{3},\ldots,\nu_{i})&-K^{F}_{i-3}(\nu_{3},\ldots,\nu_{i-1})\end{matrix}\right),\quad i=2,\ldots,r.$
Lemma 5.
$K^{F}_{r-1}(\nu_{2},\ldots,\nu_{r})=1.$
Proof.
Denoting by $\left(\begin{smallmatrix}K^{\prime}&\ast\\ -K^{\prime\prime}&\ast\end{smallmatrix}\right)$ the matrix given by (6) for $i=r$, we have
$a_{r+1}=K^{\prime}a_{2}-K^{\prime\prime}a_{1}.$
Taking stock of (5) and using the fact that $a_{1}/q_{1}<a_{2}/q_{2}$ are consecutive elements in $\mathcal{F}_{Q}$, so $a_{2}q_{1}-a_{1}q_{2}=1$, we find
$1=a_{r+1}q_{1}-a_{1}q_{r+1}=K^{\prime}(a_{2}q_{1}-a_{1}q_{2})=K^{\prime},$
as desired.
Lemma 6.
For every $r\geq 2$, we have $q_{1}+q_{r+1}\leq Q$.
Proof.
Suppose, by contradiction, that $q_{1}+q_{r+1}>Q$. Equality (5) yields $\gcd(q_{1},q_{r+1})=1$. Using the inequality $q_{1}+q_{r+1}>Q$ and the bijection between pairs of denominators of consecutive elements in $\mathcal{F}_{Q}$ and primitive lattice points $(q,q^{\prime})$ in the triangle $Q\mathcal{T}$, it follows that $q_{1}$ and $q_{r+1}$ are denominators of two consecutive elements in $\mathcal{F}_{Q}$. But this is a contradiction because $r\geq 2$ and $a_{1}/q_{1}<a_{2}/q_{2}<a_{r+1}/q_{r+1}$. ∎
Lemma 7.
For every $r\geq 2$, we have $3q_{1}+2q_{r+1}>Q$.
Proof.
Upon Lemma 6, $q_{1}+q_{r+1}=q_{i}$ for some $i\in\{2,\ldots,r\}$. We have
$(q_{i+1},q_{i})=(q_{2},q_{1})\left(\begin{matrix}a&b\\ c&d\end{matrix}\right)=(q_{r+1},q_{r})\left(\begin{matrix}\alpha&\beta\\ \gamma&\delta\end{matrix}\right)$
for some integers $a,b,c,d,\alpha,\beta,\gamma,\delta$ with $ad-bc=1=\alpha\delta-\beta\gamma$. Lemma 5 yields
$(q_{r+1},q_{r})=(q_{2},q_{1})\left(\begin{matrix}1&\nu\\ -K&1-K\nu\end{matrix}\right)$ (7)
for some (positive) integers $K,\nu$, showing that $q_{r}=q_{1}+\nu q_{r+1}$. Hence, we gather
$\begin{cases}q_{i+1}=\alpha q_{r+1}+\gamma q_{r}=\alpha q_{r+1}+\gamma(q_{1}+\nu q_{r+1})=\gamma q_{1}+(\alpha+\nu\gamma)q_{r+1}\\ q_{i}=\beta q_{r+1}+\delta q_{r}=\beta q_{r+1}+\delta(q_{1}+\nu q_{r+1})=\delta q_{1}+(\beta+\nu\delta)q_{r+1},\end{cases}$
which leads to
$(q_{i+1},q_{i})=(q_{1},q_{r+1})\left(\begin{matrix}A&B\\ C&D\end{matrix}\right),$
with $A=\gamma$, $B=\delta$, $C=\alpha+\nu\gamma$, $D=\beta+\nu\delta$ satisfying $AD-BC=\beta\gamma-\alpha\delta=-1$. Since $q_{i}=q_{1}+q_{r+1}$, we find $B=D=1$ and so $C=A+1$ and $q_{i+1}=Aq_{1}+(A+1)q_{r+1}$ with $A\geq 0$. This yields $\lfloor q_{i+1}/q_{i}\rfloor=\lfloor(Aq_{1}+(A+1)q_{r+1})/(q_{1}+q_{r+1})\rfloor=A$.
Applying one of the inequalities in [3, (16)] we infer
$\begin{split}Q&<(2+\lfloor q_{i+1}/q_{i}\rfloor)q_{i}-q_{i+1}+a_{i}\\ &\leq(3+A)q_{i}-q_{i+1}=(3+A)(q_{1}+q_{r+1})-\big{(}Aq_{1}+(A+1)q_{r+1}\big{)}=3q_{1}+2q_{r+1},\end{split}$
as desired. ∎
Proposition 8.
For every $r\geq 2$, we have
$\mathscr{D}_{Q}^{(r)}\subseteq Q\mathcal{V}_{2}\cup Q\mathcal{V}_{3}.$
Proof.
The equality (7) yields $q_{r}=q_{1}+\nu q_{r+1}$. It is plain that $K\geq 1$ and $\nu\geq 1$ in (7). Indeed, if $K\leq-1$, then $Q\geq q_{r+1}=q_{2}-Kq_{1}\geq q_{2}+q_{1}>Q$, a contradiction, while $K=0$ yields $(q_{r+1},q_{r})=(q_{2},q_{1})$, a contradiction as $r>1$. Hence $K\geq 1$. Next, if $\nu\leq-1$, then $q_{r}\leq q_{1}-q_{r+1}$, so $Q<q_{r}+q_{r+1}\leq q_{1}\leq Q$, a contradiction. Finally, if $\nu=0$, then $q_{r}=q_{1}$ and $Q<q_{r}+q_{r+1}=q_{1}+q_{2}-Kq_{1}\leq q_{2}\leq Q$, a contradiction. Hence $\nu\geq 1$ and we infer $q_{1}+q_{r+1}\leq q_{1}+\nu q_{r+1}=q_{r}\leq Q$.
It was observed in Sections 3.2 and 3.3 of [3] that when $r\geq 2$ we have $q_{1}<q_{2}$. Next, we prove the inequality $q_{r+1}-2q_{1}+Q\geq 0$. This is obsolete when $K\geq 2$, as in this case we can use $q_{r+1}=q_{2}-Kq_{1}$ to infer $q_{1}\leq q_{2}/K\leq Q/2$. When $K=1$, we have $q_{r+1}-2q_{1}=q_{2}-3q_{1}$. Now, the first inequality in [3, (16)], where $q=q_{2}>\alpha=q_{1}$, yields $3q_{1}-q_{2}\leq(2+[q_{2}/q_{1}])q_{1}-q_{2}+a_{1}\leq Q$, which leads to $q_{2}-3q_{1}+Q\geq 0$.
Finally, the inequality $3q_{1}+2q_{r+1}\geq Q$ was proved in Lemma 5. ∎
3. The inclusion $\mathcal{V}\subseteq\overline{\mathscr{D}}$
We now show that $\mathcal{V}$ cannot be further restricted; that is that $\mathcal{V}\subseteq\overline{\mathscr{D}}$, the closure of the set $\mathscr{D}$ defined in Theorem 1. In particular, we prove the following density result.
Proposition 9.
For any $(x_{0},y_{0})\in\mathring{\mathcal{V}}_{1}\cup\mathring{\mathcal{V}}_{2}\cup\mathring{\mathcal{V}}_{3}$ and $\varepsilon_{1},\varepsilon_{2}>0$, there are $Q\geq 3$ and $q_{1},q_{2}$ such that
$(q_{1},q_{2})\in\mathscr{D}_{Q}\cap Q\big{(}[x_{0}-\varepsilon_{1},x_{0}+\varepsilon_{1}]\times[y_{0}-\varepsilon_{2},y_{0}+\varepsilon_{2}]\big{)}.$
We simply denote $\mathcal{R}:=[x_{0}-\varepsilon_{1},x_{0}+\varepsilon_{1}]\times[y_{0}-\varepsilon_{2},y_{0}+\varepsilon_{2}]=:I_{1}\times I_{2}$.
The statement of Theorem 2 will be proved in parallel.
We start off with the easiest case $x_{0}+y_{0}>1$ and assume, without loss of generality, that $\mathcal{R}\subseteq\mathring{\mathcal{V}}_{1}$. Here we can look for $q_{1},q_{2}$ denominators of consecutive elements in $\mathcal{F}_{Q}$, so that $q_{1}+q_{2}>Q$. The considerations from Section 3.1 of [3] (see especially Corollary 6 thereby) yield
  • (i)
    $h(a_{1}/q_{1})=q_{1}+a_{1}+\bar{a}_{1}=3q_{1}-q_{2}-\bar{q}_{2}$, so $a_{1}/q_{1}\in\mathscr{S}_{Q}\Longleftrightarrow 3q_{1}-q_{2}-\bar{q}_{2}\leq Q$.
  • (ii)
    $q_{2}<q_{1}$, so $h(a_{1}/q_{1})\leq Q\Longrightarrow h(a_{2}/q_{2})=q_{2}+a_{2}+\bar{\bar{a}}_{2}\leq Q$.
This shows that we need to find at least one integer triple $(q_{1},q_{2},b)$ such that [^1] Here, $b$ plays the role of $\bar{q}_{2}=q_{1}-a_{1}$.
$q_{1}\in QI_{1},\quad q_{2}\in QI_{2},\quad q_{2}b\equiv-1\left(\mathrm{mod\;}q_{1}\right),\quad 3q_{1}-q_{2}-Q<b<q_{1}.$
The inequalities $\max\{Q-q_{1},2q_{1}-Q\}<q_{2}<q_{1}$ are plain consequences of the inclusion $\mathcal{R}=I_{1}\times I_{2}\subseteq\mathring{\mathcal{V}}_{1}$. The set of such integer triples is given by
$\begin{split}\sum\limits_{q_{1}\in QI_{1}}&\bigg{(}\frac{\varphi(q_{1})}{q_{1}^{2}}\operatorname{Area}\big{\{}(q_{2},b):q_{2}\in QI_{2},3q_{1}-q_{2}-Q<b<q_{1}\big{\}}+O_{\varepsilon}(q^{1/2+\varepsilon})\bigg{)}\\ &=\frac{Q^{2}}{\zeta(2)}\int_{QI_{1}}\frac{dx}{x}\operatorname{Area}\big{\{}(y,z):y\in I_{2},3x-y-1<z<x\big{\}}+O_{\varepsilon}(Q^{3/2+\varepsilon}).\end{split}$
Since 3xy1<x3x-y-1<x on $\mathcal{R}$, this is equal to
$\frac{Q^{2}}{\zeta(2)}\iint_{\mathcal{R}}H_{1}(x,y)\,dxdy+O_{\varepsilon}(Q^{3/2+\varepsilon}),$
with
$H_{1}(x,y)=\frac{1+y-2x}{x}>0\quad\text{ on $\mathring{\mathcal{V}}_{1}$},$ (8)
allowing us to conclude the proof.
We now look at the case where $x_{0}+y_{0}<1$. Although this case is similar to the previous one, it is a bit more involved since there will be one or possibly more fractions in $\mathcal{F}_{Q}$ between $a_{1}/q_{1}$ and $a_{2}/q_{2}$. In order to deal with these intermediate fractions, we use the following result:
Lemma 10 (Lemma A.1 in [3]).
Let $a_{1}/q_{1}<a_{2}/q_{2}$, $0\leq a_{i}<q_{i}$, with $a_{2}q_{1}-a_{1}q_{2}=1$. Set $Q_{i}:=h(a_{i}/q_{i})$ and $Q_{0}:=\max\{Q_{1},Q_{2}\}$. Consider the mediant $a_{\ast}/q_{\ast}:=(a_{1}+a_{2})/(q_{1}+q_{2})$ and set $Q_{\ast}:=h(a_{\ast}/q_{\ast})$. Then
  1. $(1)$
    $Q_{0}<Q_{\ast}$.
  2. $(2)$
    If we also assume that $a_{1}/q_{1}<a_{2}/q_{2}$ are consecutive elements in $\mathscr{S}_{Q_{0}}$, then $a_{1}/q_{1}<a_{2}/q_{2}$ are also consecutive elements in $\mathscr{S}_{Q}$ if $Q_{0}\leq Q<Q_{\ast}$, while $a_{1}/q_{1}<a_{\ast}/q_{\ast}<a_{2}/q_{2}$ are consecutive in $\mathscr{S}_{Q_{\ast}}$.
Therefore,
$\begin{split}a_{1}/q_{1}<a_{2}/q_{2}\text{ are consecutive in $\mathscr{S}_{Q}$ }&\Longleftrightarrow\ a_{1}/q_{1},a_{2}/q_{2}\in\mathscr{S}_{Q}\text{ and }a_{\ast}/q_{\ast}\notin\mathscr{S}_{Q}\\ &\Longleftrightarrow\ \max\{h(a_{1}/q_{1}),h(a_{2}/q_{2})\}\leq Q<h(a_{\ast}/q_{\ast}).\end{split}$ (9)
Recall the following:
Lemma 11 (Lemma 5 in [3]).
Suppose that $a_{1}/q_{1}<a_{2}/q_{2}$, $0<a_{i}<q_{i}$, and $a_{2}q_{1}-a_{1}q_{2}=1$. Denote by $\bar{n}$ the multiplicative inverse of $n\left(\mathrm{mod\;}q_{1}\right)$ in $1,q_{1})$ and by $\bar{\bar{n}}$ the multiplicative inverse of $n\left(\mathrm{mod\;}q_{2}\right)$ in $1,q_{2})$. Then
  • $\rm(i)$
    $\bar{a}_{1}=\bigg{(}1+\bigg{\lfloor}\frac{q_{2}}{q_{1}}\bigg{\rfloor}\bigg{)}q_{1}-q_{2},\quad\bar{\bar{a}}_{2}=q_{1}-\bigg{\lfloor}\frac{q_{1}}{q_{2}}\bigg{\rfloor}q_{2},$
  • $\rm(ii)$
    $a_{1}=q_{1}-\bar{q}_{2},\quad a_{2}=\bar{\bar{q}}_{1},\quad a_{1}=\frac{q_{1}\bar{\bar{q}}_{1}-1}{q_{2}},\quad a_{2}=q_{2}-\frac{q_{2}\bar{q}_{2}-1}{q_{1}}.$
  • $\rm(iii)$
    $h(a_{1}/q_{1})=\bigg{(}2+\bigg{\lfloor}\frac{q_{2}}{q_{1}}\bigg{\rfloor}\bigg{)}q_{1}-q_{2}+\frac{q_{1}\bar{\bar{q}}_{1}-1}{q_{2}},$
  • $\rm(iv)$
    $h(a_{2}/q_{2})=q_{1}+\bigg{(}2-\bigg{\lfloor}\frac{q_{1}}{q_{2}}\bigg{\rfloor}\bigg{)}q_{2}-\frac{q_{2}\bar{q}_{2}-1}{q_{1}}.$
The following expressions of $h(a/q)$ will be helpful:
Lemma 12.
Under the assumptions of the previous lemma, we have
  • $\rm(i)$
    $h(a_{1}/q_{1})=\bigg{(}3+\bigg{\lfloor}\frac{q_{2}}{q_{1}}\bigg{\rfloor}\bigg{)}q_{1}-q_{2}-\bar{q}_{2}=\bigg{(}2+\bigg{\lfloor}\frac{q_{2}}{q_{1}}\bigg{\rfloor}\bigg{)}q_{1}-q_{2}+\frac{q_{1}\bar{\bar{q}}_{1}-1}{q_{2}}$.
  • $\rm(ii)$
    $h(a_{2}/q_{2})=\bigg{(}1-\bigg{\lfloor}\frac{q_{1}}{q_{2}}\bigg{\rfloor}\bigg{)}q_{2}+q_{1}+\bar{\bar{q}}_{1}=q_{1}+\bigg{(}2-\bigg{\lfloor}\frac{q_{1}}{q_{2}}\bigg{\rfloor}\bigg{)}q_{2}-\frac{q_{2}\bar{q}_{2}-1}{q_{1}}$.
  • $\rm(iii)$
    $h(a_{\ast}/q_{\ast})=3q_{1}+q_{2}+\bar{\bar{q}}_{1}-\bar{q}_{2}$.
Proof.
(i) and (ii) are plain consequences of [3, Lemma 5].
For (iii), we employ (i) and $a_{2}q_{\ast}-a_{\ast}q_{2}=1$ to get
$h(a_{\ast}/q_{\ast})=3q_{\ast}-q_{2}-\bar{\bar{\bar{q}}}_{2}=3q_{1}+2q_{2}-\bar{\bar{\bar{q}}}_{2}=3q_{1}+q_{2}+\bar{\bar{q}}_{1}-\bar{q}_{2},$
where $\bar{\bar{\bar{q}}}_{2}=q_{\ast}-a_{\ast}=q_{2}+\bar{q}_{2}-\bar{\bar{q}}_{1}$ denotes the multiplicative inverse of $q_{2}\left(\mathrm{mod\;}q_{\ast}\right)$ in $1,q_{\ast})$. ∎
Assume first that $(x_{0},y_{0})\in\mathring{\mathcal{V}}_{2}$, so $(x_{0},y_{0})\in\mathring{\mathcal{V}}$, $x_{0}+y_{0}<1$, and $x_{0}<y_{0}$. We prove that for large $Q$ there exist integer triples $(q_{1},q_{2},a)$ such that $q_{2}>q_{1}$, $q_{1}a\equiv 1\left(\mathrm{mod\;}q_{2}\right)$, and [^2] Here, $a$ plays the role of $\bar{\bar{q}}_{1}$.
$\begin{split}&(q_{1},q_{2})\in Q\mathcal{R},\quad 0<a<q_{2},\\ &\bigg{(}2+\bigg{\lfloor}\frac{q_{2}}{q_{1}}\bigg{\rfloor}\bigg{)}q_{1}-q_{2}+\frac{q_{1}a-1}{q_{2}}\leq Q,\quad(\Longleftrightarrow h(a_{1}/q_{1})\leq Q)\\ &q_{1}+q_{2}+a\leq Q,\quad(\Longleftrightarrow h(a_{2}/q_{2})\leq Q)\\ &2q_{1}+q_{2}+a+\frac{q_{1}a-1}{q_{2}}>Q.\quad(\Longleftrightarrow h(a_{*}/q_{*})>Q)\end{split}$ (10)
The inequalities in (10) are equivalent to
$\begin{split}F_{Q}(q_{1},q_{2})&:=\max\bigg{\{}\frac{1+q_{2}(Q-2q_{1}-q_{2})}{q_{1}+q_{2}},0\bigg{\}}<a\\ &\leq G_{Q}(q_{1},q_{2}):=\min\bigg{\{}\frac{1+q_{2}(Q+q_{2}-(2+\lfloor q_{2}/q_{1}\rfloor)q_{1})}{q_{1}},Q-q_{1}-q_{2},q_{2}\bigg{\}}.\end{split}$
By a plain check, we see that
$\begin{split}&F_{Q}(q_{1},q_{2})\leq\frac{1+q_{2}(Q-2q_{1})}{q_{1}}\leq\frac{1+q_{2}(Q+q_{2}-(2+\lfloor q_{2}/q_{1}\rfloor)q_{1})}{q_{1}},\\ &F_{Q}(q_{1},q_{2})\leq Q-q_{1}-q_{2},\quad\text{and}\\ &F_{Q}(q_{1},q_{2})\leq q_{2},\end{split}$
where for the latter we made use of the inequality 3x+2y1>03x+2y-1>0 on $\mathring{\mathcal{V}}_{2}$.
Employing standard techniques, we find that the number of triples $(q_{1},q_{2},a)$ as in (10) is given by
$\frac{Q^{2}}{\zeta(2)}\iint_{\mathcal{R}}H_{2}(x,y)\,dxdy+O_{\varepsilon}(Q^{3/2+\varepsilon}),$
where
$H_{2}(x,y)=\min\bigg{\{}\frac{1+y-(2+\lfloor y/x\rfloor)x}{x},\frac{1-x-y}{y},1\bigg{\}}-\max\bigg{\{}\frac{1-2x-y}{x+y},0\bigg{\}},$ (11)
with $H_{2}(x,y)>0$ on $\mathring{\mathcal{V}}_{2}$.
Assume next that $(x_{0},y_{0})\in\mathring{\mathcal{V}}_{3}$, so $(x_{0},y_{0})\in\mathcal{V}$, $x_{0}+y_{0}<1$ and $y_{0}<x_{0}$. We prove that for large $Q$ there exist integer triples $(q_{1},q_{2},b)$ such that $q_{2}<q_{1}$, $q_{2}b\equiv-1\left(\mathrm{mod\;}q_{1}\right)$, and [^3] Here, $b$ plays the role of $\bar{q}_{2}=q_{1}-a_{1}$.
$\begin{split}&(q_{1},q_{2})\in Q\mathcal{R},\quad 0<b<q_{1},\\ &3q_{1}-q_{2}-b\leq Q,\quad(\Longleftrightarrow h(a_{1}/q_{1})\leq Q)\\ &q_{1}+\bigg{(}2-\bigg{\lfloor}\frac{q_{1}}{q_{2}}\bigg{\rfloor}\bigg{)}q_{2}-\frac{q_{2}b-1}{q_{1}}\leq Q,\quad(\Longleftrightarrow h(a_{2}/q_{2})\leq Q)\\ &3q_{1}+2q_{2}-b-\frac{q_{2}b-1}{q_{1}}>Q.\quad(\Longleftrightarrow h(a_{*}/q_{*})>Q)\end{split}$ (12)
The inequalities in (12) are equivalent to
$\begin{split}F_{Q}(q_{1},q_{2})&:=\max\bigg{\{}3q_{1}-q_{2}-Q,\frac{1+q_{1}(q_{1}+(2-\lfloor q_{1}/q_{2}\rfloor q_{2}-Q)}{q_{2}},0\bigg{\}}\leq b\\ &<G_{Q}(q_{1},q_{2}):=\min\bigg{\{}\frac{1+q_{1}(3q_{1}+2q_{2}-Q)}{q_{1}+q_{2}},q_{1}\bigg{\}}.\end{split}$
By a plain check, we see that
$F_{Q}(q_{1},q_{2})\leq\max\bigg{\{}3q_{1}-q_{2}-Q,\frac{1+q_{1}(3q_{2}-Q)}{q_{2}},0\bigg{\}}<G_{Q}(q_{1},q_{2}).$
Employing standard techniques, we find that the number of triples $(q_{1},q_{2},a)$ as in (12) is given by
$\frac{Q^{2}}{\zeta(2)}\iint_{\mathcal{R}}H_{3}(x,y)\,dxdy+O_{\varepsilon}(Q^{3/2+\varepsilon}),$
where
$H_{3}(x,y)=\min\bigg{\{}\frac{3x+2y-1}{x+y},1\bigg{\}}-\max\bigg{\{}\frac{3x-y-1}{x},\frac{x+(2-\lfloor x/y\rfloor)y-1}{y},0\bigg{\}},$ (13)
with $H_{3}(x,y)>0$ on $\mathring{\mathcal{V}}_{3}$.
Remark 1.
Employing (9), (10) and (12) to characterize consecutive elements in $\mathscr{S}_{Q}$, one can analyze in a different way than in [3] the gap distribution in $\mathscr{S}_{Q}$.
4. A partition of the region $\mathcal{V}$
Here, we assume that the sequence of fractions $a_{1}/q_{1}<\cdots<a_{r+1}/q_{r+1}$ are consecutive elements in $\mathcal{F}_{Q}$ with $a_{1}/q_{1},a_{r+1}/q_{r+1}\in\mathscr{S}_{Q}$ and $a_{2}/q_{2},\ldots,a_{r}/q_{r}\notin\mathscr{S}_{Q}$.
$\bullet$ When $r=1$, $a_{1}/q_{1}<a_{2}/q_{2}$ are consecutive in both $\mathscr{S}_{Q}$ and $\mathcal{F}_{Q}$. Lemma 4 provides
$\bigg{(}\frac{q_{1}}{Q},\frac{q_{2}}{Q}\bigg{)}\in\mathcal{W}_{1}:=\mathcal{V}_{1}\subseteq\bigg{[}0,\frac{1}{2}\bigg{]}\times\bigg{[}\frac{1}{3},1\bigg{]}.$
$\bullet$ When $r=2$, Lemma 5 provides $K^{F}_{1}(\nu_{2})=\nu_{2}=1$, so $q_{2}=q_{1}+q_{3}$. We gather $q_{1}+q_{3}\leq Q$, $2q_{1}+q_{3}=q_{1}+q_{2}>Q$, $q_{1}+2q_{3}=q_{2}+q_{3}>Q$. Here, $q_{1}<q_{2}$, so we also have $2q_{1}-q_{3}=3q_{1}-q_{2}\leq(2+[q_{2}/q_{1}])q_{1}-q_{2}\leq Q$. These inequalities translate into
$\bigg{(}\frac{q_{1}}{Q},\frac{q_{3}}{Q}\bigg{)}\in\mathcal{W}_{2}:=\bigg{\{}(x,y)\in[0,1]^{2}:\max\bigg{\{}1-2x,\frac{1-x}{2},2x-1\bigg{\}}\leq y\leq 1-x\bigg{\}},$
with $\mathcal{W}_{2}\subseteq[0,2/3]\times[1/5,1]$.
$\bullet$ When $r=3$, Lemma 5 provides $K^{F}_{2}(\nu_{2},\nu_{3})=\nu_{2}\nu_{3}-1=1$, so $(\nu_{2},\nu_{3})\in\{(2,1),(1,2)\}$.
Case 3.1. $(\nu_{2},\nu_{3})=(2,1)$, so $q_{3}=2q_{2}-q_{1}$, $q_{4}=q_{3}-q_{2}=q_{2}-q_{1}$, providing $q_{2}=q_{1}+q_{4}$, $q_{3}=q_{1}+2q_{4}\leq Q$, $q_{1}+q_{2}=2q_{1}+q_{4}>Q$, and $q_{3}+q_{4}=q_{1}+3q_{4}>Q$. We also have $q_{1}<q_{2}$ and $2q_{1}-q_{4}=3q_{1}-q_{2}\leq(2+\lfloor q_{2}/q_{1}\rfloor)q_{1}-q_{2}\leq Q$, thus
$\bigg{(}\frac{q_{1}}{Q},\frac{q_{4}}{Q}\bigg{)}\in\mathcal{W}_{3,1}:=\bigg{\{}(x,y)\in[0,1]^{2}:\max\bigg{\{}1-2x,\frac{1-x}{3},2x-1\bigg{\}}\leq y\leq\frac{1-x}{2}\bigg{\}},$
with $\mathcal{W}_{3,1}\subseteq[1/3,3/5]\times[1/7,1/3]$.
Case 3.2. $(\nu_{2},\nu_{3})=(1,2)$, so $q_{3}=q_{2}-q_{1}$, $q_{4}=2q_{3}-q_{2}=q_{2}-2q_{1}$, providing $q_{2}=2q_{1}+q_{4}\leq Q$, $q_{3}=q_{1}+q_{4}$, $q_{1}+q_{2}=3q_{1}+q_{4}>Q$, $q_{3}+q_{4}=q_{1}+2q_{4}>Q$, thus
$\bigg{(}\frac{q_{1}}{Q},\frac{q_{4}}{Q}\bigg{)}\in\mathcal{W}_{3,2}:=\bigg{\{}(x,y)\in[0,1]^{2}:\max\bigg{\{}1-3x,\frac{1-x}{2}\bigg{\}}\leq y\leq 1-2x\bigg{\}},$
with $\mathcal{W}_{3,2}\subseteq[0,1/3]\times[1/3,1]$.
$\bullet$ When $r=4$, the equation $K^{F}_{3}(\nu_{2},\nu_{3},\nu_{4})=\nu_{2}\nu_{3}\nu_{4}-\nu_{2}-\nu_{4}=1$ has five positive integer solutions: $(1,2,2)$, $(1,3,1)$, $(2,2,1)$, $(2,1,3)$, $(3,1,2)$. The solution $(2,1,3)$ cannot occur here because $q_{3}=2q_{2}-q_{1}$, $q_{4}=q_{3}-q_{2}=q_{2}-q_{1}$, $q_{5}=3q_{4}-q_{3}=q_{2}-2q_{1}$ would lead to $Q\geq q_{3}>2q_{2}-3q_{1}=q_{4}+q_{5}>Q$, a contradiction. The solution $(3,1,2)$ cannot occur either because $q_{3}=3q_{2}-q_{1}$, $q_{4}=q_{3}-q_{2}=2q_{2}-q_{1}$, $q_{5}=2q_{4}-q_{3}=q_{2}-q_{1}$ would lead to $Q\geq q_{3}>3q_{2}-2q_{1}=q_{4}+q_{5}>Q$, a contradiction. We are left with three cases:
Case 4.1. $(\nu_{2},\nu_{3},\nu_{4})=(2,2,1)$, so $q_{3}=2q_{2}-q_{1}$, $q_{4}=2q_{3}-q_{2}=3q_{2}-2q_{1}$, $q_{5}=q_{4}-q_{3}=q_{2}-2q_{1}$, providing $q_{2}=q_{1}+q_{5}$, $q_{3}=q_{1}+2q_{5}$, $q_{4}=q_{1}+3q_{5}\leq Q$, and thus $q_{1}+q_{2}=2q_{1}+q_{5}>Q$ and $q_{4}+q_{5}=q_{1}+4q_{5}>Q$. Again, the inequality $q_{5}\geq 2q_{1}-Q$ follows for free since $q_{1}<q_{2}$ and $2q_{1}-q_{5}=3q_{1}-q_{2}\leq(2+\lfloor q_{2}/q_{1}\rfloor)q_{1}-q_{2}\leq Q$. Thus
$\bigg{(}\frac{q_{1}}{Q},\frac{q_{5}}{Q}\bigg{)}\in\mathcal{W}_{4,1}:=\bigg{\{}(x,y)\in[0,1]^{2}:\max\bigg{\{}1-2x,\frac{1-x}{4},2x-1\bigg{\}}\leq y\leq\frac{1-x}{3}\bigg{\}},$
with $\mathcal{W}_{4,1}\subseteq[2/5,4/7]\times[1/9,1/5]$.
Case 4.2. $(\nu_{2},\nu_{3},\nu_{4})=(1,3,1)$, so $q_{3}=q_{2}-q_{1}$, $q_{4}=3q_{3}-q_{2}=2q_{2}-3q_{1}$, $q_{5}=q_{4}-q_{3}=q_{2}-2q_{1}$, providing $q_{2}=2q_{1}+q_{5}\leq Q$, $q_{3}=q_{1}+q_{5}$, $q_{4}=q_{1}+2q_{5}\leq Q$. Thus $q_{1}+q_{2}=3q_{1}+q_{5}>Q$ and $q_{4}+q_{5}=q_{1}+3q_{5}>Q$, leading to
$\bigg{(}\frac{q_{1}}{Q},\frac{q_{5}}{Q}\bigg{)}\in\mathcal{W}_{4,2}:=\bigg{\{}(x,y)\in[0,1]^{2}:\max\bigg{\{}1-3x,\frac{1-x}{3}\bigg{\}}\leq y\leq\min\bigg{\{}\frac{1-x}{2},1-2x\bigg{\}}\bigg{\}},$
with $\mathcal{W}_{4,2}\subseteq[1/5,2/5]^{2}$.
Case 4.3. $(\nu_{2},\nu_{3},\nu_{4})=(1,2,2)$, so $q_{3}=q_{2}-q_{1}$, $q_{4}=2q_{3}-q_{2}=q_{2}-2q_{1}$, $q_{5}=2q_{4}-q_{3}=q_{2}-3q_{1}$, providing $q_{2}=3q_{1}+q_{5}\leq Q$, $q_{3}=2q_{1}+q_{5}$, $q_{4}=q_{1}+q_{5}$. Thus $q_{1}+q_{2}=3q_{1}+q_{5}>Q$ and $q_{4}+q_{5}=q_{1}+2q_{5}>Q$, leading to
$\bigg{(}\frac{q_{1}}{Q},\frac{q_{5}}{Q}\bigg{)}\in\mathcal{W}_{4,3}:=\bigg{\{}(x,y)\in[0,1]^{2}:\max\bigg{\{}1-4x,\frac{1-x}{2}\bigg{\}}\leq y\leq 1-3x\bigg{\}},$
with $\mathcal{W}_{4,3}\subseteq[0,1/5]\times[2/5,1]$.
$\bullet$ When $r=5$, the equation $K^{F}_{4}(\nu_{2},\nu_{3},\nu_{4},\nu_{5})=\nu_{2}\nu_{3}\nu_{4}\nu_{5}-\nu_{2}\nu_{3}-\nu_{2}\nu_{5}-\nu_{4}\nu_{5}+1=1$ has twelve positive solutions:$(1,3,1,3)$, $(1,4,1,2)$, $(2,1,3,2)$, $(2,1,4,1)$, $(2,2,1,4)$, $(3,1,2,3)$, $(3,1,3,1)$, $(3,2,1,3)$, $(2,2,2,1)$, $(1,3,2,1)$, $(1,2,3,1)$, $(1,2,2,2)$. The first eight solutions are not admissible because
$\bullet$ $(1,3,1,3)$ would yield $Q\geq q_{4}=2q_{2}-3q_{1}>2q_{2}-5q_{1}=q_{5}+q_{6}>Q$, a contradiction.
$\bullet$ $(1,4,1,2)$ would yield $Q\geq q_{4}=3q_{2}-4q_{1}>3q_{2}-5q_{1}=q_{5}+q_{6}>Q$, a contradiction.
$\bullet$ $(2,1,3,2)$ would yield $Q\geq q_{3}=2q_{2}-q_{1}>2q_{2}-3q_{1}=q_{4}+q_{5}>Q$, a contradiction.
$\bullet$ $(2,1,4,1)$ would yield $q_{3}=2q_{2}-q_{1}$, $q_{4}=q_{3}-q_{2}=q_{2}-q_{1}$, $q_{5}=2q_{2}-3q_{1}$, and $q_{6}=q_{5}-q_{4}=q_{2}-2q_{1}>0$, so $q_{2}>2q_{1}$, which gives $\nu_{2}=1$, a contradiction.
$\bullet$ $(2,2,1,4)$ would yield $Q\geq q_{3}=2q_{2}-q_{1}>2q_{2}-3q_{1}=q_{5}+q_{6}>Q$, a contradiction.
$\bullet$ $(3,1,2,3)$ would yield $Q\geq q_{4}=2q_{2}-q_{1}>2q_{2}-3q_{1}=q_{5}+q_{2}>Q$, a contradiction.
$\bullet$ $(3,1,3,1)$ would yield $q_{2}>q_{1}$, so $\nu_{2}\in\{1,2\}$, a contradiction.
$\bullet$ $(3,2,1,3)$ would yield $Q\geq q_{3}=3q_{2}-q_{1}>3q_{2}-2q_{1}=q_{5}+q_{6}>Q$, a contradiction.
We are left with four cases:
Case 5.1. $(\nu_{2},\nu_{3},\nu_{4},\nu_{5})=(2,2,2,1)$, so $q_{3}=2q_{2}-q_{1}$, $q-4=2q_{3}-q_{2}=3q_{2}-2q_{1}$ $q_{5}=2q_{4}-q_{3}$, $q_{6}=q_{5}-q_{4}=q_{2}-q_{1}$, providing $q_{2}=q_{1}+q_{6}$, $q_{3}=q_{1}+2q_{6}$, $q_{4}=q_{1}+3q_{6}$, $q_{5}=q_{1}+4q_{6}\leq Q$. Thus $q_{1}+q_{2}=2q_{1}+q_{6}>Q$ and $q_{1}+5q_{6}=q_{5}+q_{6}>Q$, leading to
$\bigg{(}\frac{q_{1}}{Q},\frac{q_{6}}{Q}\bigg{)}\in\mathcal{W}_{5,1}:=\bigg{\{}(x,y)\in[0,1]^{2}:\max\bigg{\{}1-2x,\frac{1-x}{5},2x-1\bigg{\}}\leq y\leq\frac{1-x}{4}\bigg{\}},$
with $\mathcal{W}_{5,1}\subseteq[3/7,6/11]\times[1/11,1/7]$.
Case 5.2. $(\nu_{2},\nu_{3},\nu_{4},\nu_{5})=(1,3,2,1)$, so $q_{3}=q_{2}-q_{1}$, $q_{4}=3q_{3}-q_{2}=2q_{2}-3q_{1}$, $q_{5}=2q_{4}-q_{3}=3q_{2}-5q_{1}$, $q_{6}=q_{5}-q_{4}=q_{2}-2q_{1}$, providing $q_{2}=2q_{1}+q_{6}\leq Q$, $q_{3}=q_{1}+q_{6}$, $q_{4}=q_{1}+2q_{6}$, $q_{5}=q_{1}+3q_{6}\leq Q$. Thus $q_{1}+q_{2}=2q_{1}+q_{6}>Q$ and $q_{5}+q_{6}=q_{1}+4q_{6}>Q$, leading to
$\bigg{(}\frac{q_{1}}{Q},\frac{q_{6}}{Q}\bigg{)}\in\mathcal{W}_{5,2}:=\bigg{\{}(x,y)\in[0,1]^{2}:\max\bigg{\{}1-3x,\frac{1-x}{4}\bigg{\}}\leq y\leq\min\bigg{\{}\frac{1-x}{3},1-2x\bigg{\}}\bigg{\}},$
with $\mathcal{W}_{5,2}\subseteq[1/4,3/7]\times[1/7,1/4]$.
Case 5.3. $(\nu_{2},\nu_{3},\nu_{4},\nu_{5})=(1,2,3,1)$, so $q_{3}=q_{2}-q_{1}$, $q_{4}=2q_{3}-q_{2}=q_{2}-2q_{1}$, $q_{5}=3q_{4}-q_{3}=2q_{2}-5q_{1}$, $q_{6}=q_{5}-q_{4}=q_{2}-3q_{1}$, providing $q_{2}=3q_{1}+q_{6}\leq Q$, $q_{3}=2q_{1}+q_{6}$, $q_{4}=q_{1}+q_{6}$, $q_{5}=q_{1}+2q_{6}\leq Q$. Thus $q_{1}+q_{2}=4q_{1}+q_{6}>Q$ and $q_{5}+q_{6}=q_{1}+3q_{6}>Q$, leading to
$\bigg{(}\frac{q_{1}}{Q},\frac{q_{6}}{Q}\bigg{)}\in\mathcal{W}_{5,3}:=\bigg{\{}(x,y)\in[0,1]^{2}:\max\bigg{\{}1-4x,\frac{1-x}{3}\bigg{\}}\leq y\leq\min\bigg{\{}\frac{1-x}{2},1-3x\bigg{\}}\bigg{\}},$
with $\mathcal{W}_{5,3}\subseteq[1/7,1/4]\times[1/4,3/7]$.
Case 5.4. $(\nu_{2},\nu_{3},\nu_{4},\nu_{5})=(1,2,2,2)$, so $q_{3}=q_{2}-q_{1}$, $q_{4}=2q_{3}-q_{2}=q_{2}-2q_{1}$, $q_{5}=2q_{4}-q_{3}=q_{2}-3q_{1}$. $q_{6}=2q_{5}-q_{4}=q_{2}-4q_{1}$, providing $q_{2}=4q_{1}+q_{6}\leq Q$, $q_{3}=3q_{1}+q_{6}$, $q_{4}=2q_{1}+q_{6}$, $q_{5}=q_{1}+q_{6}$. Thus $q_{1}+q_{2}=5q_{1}+q_{6}>Q$ and $q_{5}+q_{6}=q_{1}+2q_{6}>Q$, leading to
$\bigg{(}\frac{q_{1}}{Q},\frac{q_{6}}{Q}\bigg{)}\in\mathcal{W}_{5,4}:=\bigg{\{}(x,y)\in[0,1]^{2}:\max\bigg{\{}1-5x,\frac{1-x}{2}\bigg{\}}\leq y\leq 1-4x\bigg{\}},$
with $\mathcal{W}_{5,4}\subseteq[0,1/7]\times[3/7,1]$.
Next, we show the existence of examples where the boundary line $3x+2y=1$ of $\mathcal{V}$ occurs as a side of a region $\mathcal{W}_{r,i}$. The tuple $(q_{1},\ldots,q_{10})=(27,95,68,41,96,55,69,83,97,14)$ represents consecutive elements in $\mathcal{F}_{100}$ with $q_{1}=27$ and $q_{10}=84$ consecutive in $\mathscr{S}_{100}$ and mediant $q_{1}+q_{10}=41=q_{4}$. The equalities $q_{2}=3q_{1}+q_{10}$, $q_{3}=2q_{1}+q_{10}$, $q_{4}=q_{1}+q_{10}$, $q_{5}=2q_{1}+3q_{10}$, $q_{6}=q_{1}+2q_{10}$, $q_{7}=q_{1}+3q_{10}$, $q_{8}=q_{1}+4q_{10}$, $q_{9}=q_{1}+5q_{10}$ show that the region $Q\mathcal{W}_{9,4}$ is defined by the inequalities $\max\{3q_{1}+q_{10},2q_{1}+3q_{10},q_{1}+5q_{10}\}\leq Q<\min\{q_{1}+q_{2}=4q_{1}+q_{10},q_{3}+q_{4}=3q_{1}+2q_{10},q_{6}+q_{7}=2q_{1}+5q_{10},q_{9}+q_{10}=q_{1}+6q_{10}\}$, and thus
$\mathcal{W}_{9,4}=\{h_{1}(x)\leq y\leq h_{2}(x)\},$
with
$\begin{split}&h_{1}(x)=\max\bigg{\{}1-4x,\frac{1-3x}{2},\frac{1-2x}{5},\frac{1-x}{6}\bigg{\}}=\begin{cases}1-4x&\mbox{\rm if $0\leq x\leq 1/5$}\\ (1-3x)/2&\mbox{\rm if $1/5\leq x\leq 1/4$}\\ (1-x)/6&\mbox{\rm if $1/4\leq x\leq 1/2$}\end{cases},\\ &h_{2}(x)=\min\bigg{\{}1-3x,\frac{1-2x}{3},\frac{1-x}{5}\bigg{\}}=\begin{cases}(1-x)/5&\mbox{\rm if $0\leq x\leq 2/7$}\\ 1-3x&\mbox{\rm if $2/7\leq x\leq 1/2$}\end{cases}.\end{split}$
We have $\mathcal{W}_{9,4}\subseteq[3/13,5/17]\times[2/17,2/13]$ and the line $3x+2y=1$ is a side of $\mathcal{W}_{9,4}$ because
$\mathcal{W}_{9,4}\cap\bigg{(}\bigg{[}\frac{1}{5},\frac{1}{4}\bigg{]}\times[0,1]\bigg{)}=\bigg{\{}(x,y):\frac{1}{5}\leq x\leq\frac{1}{4},\frac{1-3x}{2}\leq y\leq\frac{1-x}{5}\bigg{\}}.$
5. An index identity and the $\Phi$ function
As $a_{i+1}q_{i}-a_{i}q_{i+1}=1$, we have
$\nu(\gamma_{i}):=\frac{q_{i-1}+q_{i+1}}{q_{i}}=\frac{a_{i-1}+a_{i+1}}{a_{i}}=a_{i+2}q_{i}-a_{i}q_{i+2}\in\mathbb{N},\qquad i=1,\ldots,S_{Q}.$
The proof of Proposition 3 follows closely the idea of [10, Theorem 1] and proceeds by induction on $Q$. Denote the sum from (4) by $T_{Q}$. For $Q=3$, we have $\mathscr{S}_{3}=\{1/1\}$, $S_{3}=1$, and $T_{3}=2=3R_{3}-1$. Suppose that $T_{Q-1}=3S_{Q-1}-1$ for some $Q\geq 4$. The number of mediant insertions when enlarging $\mathscr{S}_{Q-1}$ to $\mathscr{S}_{Q}$ is precisely $S_{Q}-S_{Q-1}$. When such a mediant insertion $\gamma_{\ast}=a_{\ast}/q_{\ast}$ occurs, say between $\gamma_{i}$ and $\gamma_{i+1}$, its contribution to $T_{Q-1}$ is
$A_{i}=\frac{q_{i-1}+q_{i+1}}{q_{i}}+\frac{q_{i}+q_{i+2}}{q_{i+1}},$
while its contribution to $T_{Q}$ is
$\begin{split}B_{i}&=\frac{q_{i-1}+q_{*}}{q_{i}}+\frac{q_{i}+q_{i+1}}{q_{*}}+\frac{q_{*}+q_{i+2}}{q_{i+1}}\\ &=\frac{q_{i-1}+(q_{i}+q_{i+1})}{q_{i}}+1+\frac{(q_{i}+q_{i+1})+q_{i+2}}{q_{i+1}}=A_{i}+3.\end{split}$
This shows that $T_{Q}=T_{Q-1}+3(S_{Q}-S_{Q-1})=3S_{Q}-1$, as desired.
Figure 4. The horizontal and vertical transformations when insertions are made. Thus, a point $(q_{1},q_{2})\in\mathcal{D}_{Q}$ that disappears at level $Q$ is replaced in $\mathcal{D}_{Q+1}$ with two new points $(q_{1}+q_{2},q_{2})$ and $(q_{1},q_{1}+q_{2})$. Therefore, through insertions, the polygon $ABEF$, in which the points that will disappear at level $Q$ are found, transforms into the polygons $ABCD$ and $GDEF$ that will contain the new points.
It also suggests that an analogue of Euler’s function $\varphi(Q)$ is given here by
$\begin{split}\Phi(Q)&:=S_{Q}-S_{Q-1}=\#\big{\{}a/q\in\mathbb{Q}\cap(0,1]:q+a+\bar{a}=Q\big{\}}\\ &=\#\bigg{\{}\frac{a_{1}}{q_{1}}<\frac{a_{2}}{q_{2}}\text{ consecutive in $\mathscr{S}_{Q-1}$:}\ h\bigg{(}\frac{a_{1}+a_{2}}{q_{1}+q_{2}}\bigg{)}=Q\bigg{\}}\\ &=\#\left\{(a_{1},q_{1},a_{2},q_{2})\in\mathbb{N}^{4}:\begin{matrix}a_{2}q_{1}-a_{1}q_{2}=1,\ a_{i}<q_{i},\\ \max\big{\{}h\big{(}\frac{a_{1}}{q_{1}}\big{)},h\big{(}\frac{a_{2}}{q_{2}}\big{)}\big{\}}\leq Q-1,\ h\big{(}\frac{a_{1}+a_{2}}{q_{1}+q_{2}}\big{)}=Q\end{matrix}\right\}.\end{split}$
The arithmetic function $\Phi$ has an erratic behavior (see Figure 3). It is not multiplicative and
$\begin{split}&\Phi(3)=S_{3}=1,\quad\Phi(4)=1,\quad\Phi(5)=1,\quad\Phi(6)=1,\quad\Phi(7)=2,\quad\Phi(8)=1,\\ &\Phi(9)=1,\quad\Phi(10)=4,\quad\Phi(11)=1,\quad\Phi(12)=1,\quad\Phi(13)=4,\quad\Phi(14)=2,\\ &\Phi(15)=3,\quad\Phi(16)=4,\quad\Phi(17)=1,\quad\Phi(18)=4,\quad\Phi(19)=4,\quad\Phi(20)=5.\end{split}$
Refer to caption Refer to caption
Figure 5. The points that disappear and the new points created by insertion when passing from level Q1Q-1 to level $Q$ for $Q=721,722,723$, and $724$.
In Figures 3, 4, and 5, it can be observed that not only does $\Phi(Q)$ exhibit significant oscillation, but also that at times, particularly around levels $Q$ with a large number of divisors, the disappearing and newly created points tend to group together into distinct formations.